OpenStudy (anonymous):

How many milliliters of 17% alcohol solution must be mixed with 40 milliliters of a 3% alcohol to obtain a 12% alcohol solution?

OpenStudy (xapproachesinfinity):

let x be the milliliters of 17% alcohol solution let y be the total milliliters of the resulting alcohol solution then we can say that x+40=y and 0.17x+0.03*40=0.12y

OpenStudy (xapproachesinfinity):

solve for x by replacing y=x+40 in the second equation and you are done